OpenStudy (bloomlocke367):

I forgot what to do... Simplify: \(\Huge\frac{x^2-4}{x^2+5x+6}\) State any restrictions.

OpenStudy (mathmate):

Factorize both denominator and numerator. If there are common factors, you can cancel on condition that you state that the factor cannot have a value of zero. Example: (x^2-1)/(x+1) = (x+1)(x-1)/(x-1) =x+1 on condition that x\(\ne\)1

OpenStudy (bloomlocke367):

so I have \(\huge\frac{(x+2)(x-2)}{(x+2)(x+3)}\) so it's \(\huge\frac{x-2}{x+3}\) right?

OpenStudy (bloomlocke367):

@mathmate

OpenStudy (bloomlocke367):

how do I know the restrictions?

OpenStudy (bloomlocke367):

@rational

OpenStudy (mathmate):

Before you cancel the common factor (x+2), the restriction is that the factor cannot be zero.

OpenStudy (mathmate):

\(x+2 \ne 0\) also means x\(\ne\) -2. This is the restriction.

OpenStudy (bloomlocke367):

but there's two restrictions... is the other one -3?

OpenStudy (mathmate):

Some people would state the second restriction of x+3\(\ne\)0. (i.e. the other factor of the denominator). I usually don't because it is an intrinsic restriction of the expression, not one arising from the simplification.

OpenStudy (mathmate):

Note: -3 is not a restriction, x\(\ne\)-3 is.

OpenStudy (bloomlocke367):

yes, I know that it's \(x\cancel=-3\)

OpenStudy (bloomlocke367):

thank you